### v2

# ViewDelta: Text-Prompted Change Detection in Unaligned Images

Abstract
Detecting changes between images is fundamental in applications such as infrastructure assessment, environmental monitoring, and industrial automation. Existing supervised models demonstrate strong performance but are inherently limited by the scope of their training data, requiring retraining to recognize novel changes. To overcome this limitation, we introduce a novel change detection task utilizing textual prompts alongside two potentially unaligned images to produce binary segmentations highlighting user-relevant changes. This text-conditioned framework significantly broadens the scope of change detection, enabling unparalleled flexibility and straightforward scalability by incorporating diverse future datasets without restriction to specific change types. As a first approach to address this challenge, we propose ViewDelta, a multimodal architecture extending the vision transformer into the domain of text-conditioned change detection. ViewDelta establishes a robust baseline, demonstrating flexibility across various scenarios and achieving competitive results compared to specialized, fine-tuned models trained on aligned images. Moreover, we create and release the first text-prompt-conditioned change detection dataset, comprising 501,153 image pairs with corresponding textual prompts and annotated labels. Extensive experiments confirm the robustness and versatility of our model across diverse environments, including indoor, outdoor, street-level, synthetic, and satellite imagery. github.io/viewdelta/
1 Introduction
Figure 1: Given a pair of aligned or unaligned images captured before and after an event, alongside a textual prompt specifying user-relevant changes, ViewDelta outputs a segmentation highlighting the desired changes.
Detecting and understanding changes between images is a critical task in computer vision [3, 39, 20, 9], with applications spanning situational awareness [14, 47], infrastructure damage assessment [35, 41, 45, 38], and environmental monitoring [18, 28].
Traditional change detection methods typically require pairs of aligned images and focus on predefined types of changes, limiting their applicability to scenarios where alignment cannot be guaranteed or when encountering novel change types not represented during training. Such methods thus struggle in real-world settings, which often involve varying viewpoints or camera orientations, resulting in significant performance degradation.
To overcome these limitations, we introduce ViewDelta (see Fig. 1), a novel approach enabling robust, user-defined, prompt conditioned change detection for unaligned image pairs. By leveraging natural language prompts, ViewDelta provides users the flexibility to specify desired changes explicitly, accommodating diverse applications within a single generalized model. Due to the added flexibility of a input prompt, our method can effectively handle images from a wide range of domains, such as street-level and satellite imagery, without the necessity for domain specific retraining.
Supporting this approach, we introduce CSeg, a new dataset tailored explicitly for the prompt conditoned change detection tasks involving unaligned imagery. CSeg employs advanced inpainting techniques to simulate realistic change scenarios, significantly reducing the manual labeling effort typically required. The dataset also incorporates intentional “red herring” changes, preventing the model from erroneously identifying inpainting artifacts as meaningful changes, thereby improving its generalization capability in real-world contexts.
Our extensive experiments illustrate that ViewDelta achieves competitive performance relative to specialized, domain-specific models trained exclusively on aligned images. These results hold consistently across varied settings, including indoor, outdoor, street-level, synthetic, and satellite imagery. Our contributions include:
  • Introducing a novel text prompt conditioned change detection task that outputs binary segmentations based on user-specified textual descriptions of relevant changes.
  • Proposing ViewDelta, the first transformer-based multimodal architecture explicitly designed for prompt conditioned change detection.
  • Creating and publicly releasing the CSeg dataset, comprising aligned and unaligned image pairs with corresponding textual prompts and annotated segmentation masks, facilitating future research.
  • Demonstrating the first generalized change detection model capable of robustly performing across diverse image domains without retraining.
2 Related Work
Change Detection: Change detection (CD) entails identifying variations between two images, $I_{a}$ and $I_{b}$, captured at different times and potentially from different viewpoints, producing a binary change map $M\in\{0,1\}^{H\times W}$ that indicates altered regions. Traditional CD methods have largely focused on 2D image pairs with minimal or no viewpoint changes, as commonly seen in applications such as surveillance and satellite imagery [51, 16, 17, 2]. More recent approaches have extended CD to complex 3D settings, where the camera poses of $I_{a}$ and $I_{b}$ may vary significantly, and even to inherently 3D data like point clouds [34, 43, 1, 25]. This shift toward viewpoint variability introduces additional challenges, particularly in accurately annotating changes under varying viewpoints, which makes manual annotation costly and complex.
To address the difficulties associated with labeling under variable viewpoints, some methods leverage synthetic data or 3D representations. However, these advancements have been constrained by a lack of large-scale, real-image datasets annotated for CD with viewpoint variation. Consequently, models trained in these scenarios are often limited to synthetic data or rely on 3D modalities restricting the potential use cases to only scenarios where 3D data is available. This creates a need for real-image CD models robust enough to handle diverse viewpoints. To address this gap, we leverage three datasets. For the first dataset, we use the PSCD dataset [36] which consists of shifted panoramic images that, when cropped, shows slight variation in views. For the second dataset, we use the SYSU-CD [37] dataset which consists of fixed view satellite imagery. Finally to expand the models vocabulary and introduce more variation to view, we create CSeg, a dataset consisting of multi-view real images with multiple inpainted changes.
Semantic Change Detection: Unlike traditional CD, Semantic Change Detection consists of multiple categories of changes that must be selected similarly to semantic segmentation [46, 12, 40]. Given $I_{a}$ and $I_{b}$, the objective of semantic change detection is to predict a semantic segmentation map $S\in\{1,2,\dots,N\}^{H\times W}$, where $N$ is the total number of semantic classes. Each pixel $S_{i,j}$ is assigned a class label $c\in\{1,2,\dots,N\}$, identifying the type of semantic change corresponding to that region. Semantic change detection has been successfully applied to various types of imagery, including satellite [46], street view [36], and warehouse [31] images. However, current methods are typically tailored to specific domains, and a unified model that can generalize across different categories within a single saved checkpoint does not yet exist. With ViewDelta due to our unique formulation of CD with a text prompt, we are able to jointly train on multiple datasets and create a single saved model that achieve state of the art performance rather than a finetuned model for each dataset.
Open Vocabulary Image Segmentation: Open Vocabulary Image Segmentation is a computer vision task that involves labeling each pixel with categories beyond a predefined set, including categories potentially unseen during training [26, 27, 50]. This flexibility is achieved through models that associate text and image features, often utilizing vision-language frameworks like CLIP [32], which align visual and textual embeddings within a shared space. Our approach builds on these methods by integrating text and image models to enhance change detection. Rather than highlighting all possible changes, our model interprets a user-provided text prompt to selectively highlight changes relevant to the specified user text.
3 ViewDelta
Figure 2: Overview diagram of the ViewDelta architecture consisting of two primary branches to process images and text. Both the text and image embedding network are frozen during training and never optimized. Text and image embeddings are concatenated with segmentation tokens before projection with an MLP before feeding into a ViT-Base backbone. Four layers of convolutions are applied with up sampling to decode the output features into a binary mask.
In this section, we introduce ViewDelta, our proposed architecture for text-prompted change detection, as illustrated in Fig. 2. ViewDelta processes two input images, $I_{a}$ and $I_{b}$, captured before and after a change event, respectively, along with a text prompt $T$ that specifies the type of change relevant to the user. The goal is to predict a binary segmentation map $M\in\{0,1\}^{H\times W}$, where $M_{i,j}=1$ indicates the presence of the specified change at pixel location $(i,j)$ and $M_{i,j}=0$ otherwise.
The proposed architecture introduces several key modifications to the Vision Transformer (ViT) [13] to effectively integrate text prompts. These modifications include integrating a frozen text encoder to transform the text prompt $T$ into a sequence of tokens, using an image encoder instead of a traditional patch embedding network for processing the input images $I_{a}$ and $I_{b}$, introducing learned segmentation query tokens, and employing a fully convolutional segmentation head. The segmentation head specifically produces the final binary segmentation map $M$ from the learned segmentation query tokens. These design choices primarily enable efficient training by leveraging pretrained networks, crucial given the limited number of training examples, while still exploiting the flexible but data-intensive transformer architecture.
3.1 Text Embeddings
We leverage only the text encoder of the SigLip [48] model due to its improved performance against models such as CLIP [32], OpenCLIP [15], and EVA-CLIP [42]. We leave the SigLip model frozen in order to keep the text generalization ability of SigLip. We evaluate the generalizability of the model on the CSeg dataset which contains unique prompts in the test set that are not seen in the training set.
3.2 Image Embeddings
Patch embeddings have been widely adopted by state-of-the-art methods in both change detection and semantic segmentation. However, in our joint training configuration, we observed that optimizing patch embeddings was challenging due to the slow convergence rate on the available training data. To address this issue, we investigated directly using image features from a frozen Dinov2 [29] model as embeddings. Our ablation study shall demonstrate that Dinov2 embeddings yielded faster convergence rate with less data. We attribute the difficulty of optimizing learned patch embeddings primarily to dataset size limitations rather than inherent issues with patch embeddings themselves. Scaling the model and dataset could further improve performance but would require significantly increased computational resources and data availability, as such this will be investigated further in follow up works that will scale our dataset and model size.
3.3 Backbone
We utilize the ViT [13] transformer architecture in it’s base configuration as the backbone, consisting of 12 transformer encoder layers, each characterized by a hidden dimension of 768, an intermediate MLP dimension of 3072, and 12 attention heads. Diverging from typical approaches, this backbone abstains from assumptions regarding spatial alignment and refrains from explicit feature aggregation strategies, such as subtraction or concatenation, commonly employed in models like FC-EF [11], STANet [8], [7], and DSAMNet [37]. Consequently, it achieves a generalized and robust multimodal representation without relying on a built in bias for aligned images.
3.4 Segmentation Query Tokens
Drawing inspiration from DETR [5], we leverage dedicated segmentation query tokens to enhance model flexibility. These tokens facilitate the model to create streamlined representations for segmentation-related features, circumventing the inherent complexities encountered if the model had to directly map and align combined image and textual tokens to a feature space that could be used of change detection. This methodological choice effectively addresses issues of adding padding tokens or deciding which tokens from the sequence of image and text tokens to use as input into a segmentation head. State of the art change detection methods that use transformers are able to circumvent this because images are aligned and text tokens are not incorporated.
3.5 Segmentation Head
Existing state-of-the-art change detection architectures, including SwinSUNet [49], ChangeFormerV4 [4], and MambaBCD-Base [7], predominantly assume spatial alignment between input image pairs, inherently limiting their applicability in practical scenarios involving unaligned or varying viewpoints. To overcome this limitation, we adopt a segmentation that exclusively operates only on segmentation query tokens, as shown in Algorithm 1. This design choice removes the necessity of explicit feature combination from spatially mismatched input images and facilitating effective fusion of visual and textual embeddings. The implemented segmentation head comprises of a concise structure of four convolutional layers followed by upsampling, forming a robust yet straightforward framework for investigating performance and generalization in text-conditioned change detection.
Algorithm 1 Forward prediction of proposed model
1:Images $I_{a}$, $I_{b}$ and text embeddings $T_{e}$
2:$I_{a}\leftarrow\text{ImageEmbedder}(I_{a})$
3:$I_{b}\leftarrow\text{ImageEmbedder}(I_{b})$
4:$T\leftarrow\text{TextEmbedder}(T_{e})$
5:$X\leftarrow\text{Concat}(I_{a},I_{b},T,\text{SegEmbeddings})$
[…]
7:$X\leftarrow\text{MLP}(X)$
8:$X\leftarrow\text{ViT}(X)$
[…]
11:return SegLogits
3.6 Implementation Details
PSCD view variation: The PSCD dataset comprises of 770 panoramic image pairs captured in a suburban environment. Though the PSCD dataset was originally released for view aligned change detection of panoramic images, from further inspection, we have found that, when zoomed in, the dataset was found to have some varying amounts of viewpoint change. To leverage this data as a source of real street level change detection we take each panoramic image and extract non-overlapping crops of size $256\times 256$. The labels provided by PSCD are originally intended for semantic segmentation across classes such as vehicles and people. We use these classes to generate several binary segmentation masks, with each class serving as a distinct text prompt for segmentation.
Training Details: ViewDelta is jointly trained across multiple datasets using the Adam optimizer [21] with weight decay, a one-epoch warm-up, and cosine annealing learning rate decay, starting from an initial learning rate of $2\times 10^{-5}$. Training is performed using a batch size of 4 per GPU using 4 Nvidia A100 GPUs, resulting in an effective batch size of 16. We utilize DeepSpeed ZeRO Stage 2 [33] for sharding optimizer states and gradients, activation checkpointing within attention layers and large linear layer, and 16-bit mixed-precision training.
4 CSeg Dataset
Figure 3: Overview of CSeg generation. We select an SA-1B image and leverage LLaVa-Next to propose multiple class proposals that exist in the image. These classes are then converted to a natural language change prompt and also are fed into Grounded SAM 2 to collect instance masks on all class proposals. These masks are collected and inpainted to simulate change while the SA-1B provided masks are used as red herring masks. This results in a text prompt, image A, image B, and change mask label.
A major factor contributing to recent advancements in deep learning is the accessibility of large-scale datasets with high-quality annotations. To the best of our knowledge, there is currently no existing dataset that contains image pairs with text-guided change masks, especially those that involve view shifts. Although the processed PSCD dataset provides minor view alignment differences, it is limited to traffic-like classes such as “vehicles”, “barrier”, “human”, and “lane marking”. A dataset that addresses these issues will significantly accelerate progress in the change detection field. However, manually labeling thousands of image pairs with changes is cumbersome, expensive, and impractical to scale. Inspired by the COCO-Inpainted dataset [34], we propose a procedure that leverages state of the art vision models to generate the Change Segmentation (CSeg) dataset: a large-scale, varying-view, and text-guided change detection dataset.
4.1 Inpainting Changes
Instead of manually creating and labeling change images, we employ inpainting to synthetically simulate changes between two images. To achieve this, we first select an SA-1B [22] image and leverage LLaVa-Next [24] to identify a maximum of ten different objects in the image to be used as classes. To promote dataset balance, we select one to five of the least represented classes during generation. These classes are then used directly as text prompts or applied to a natural language prompt template. These prompt templates are a list of 45 predefined and manually validated templates generated with GPT-4o that have the intent of identifying change. To simulate change, Grounded SAM 2 is then used to generate instance masks across classes, and we randomly select a subset of at most 10 masks to be inpainted in image B with LaMa [44]. We combine these inpainted masks to produce the change label. Finally, we apply a random homography change to either image A or image B to simulate a perspective change.
ViewDelta should be generalized to detect all changes in an image and not limited to those specified by a prompt which eliminates the need for explicit class enumeration. To generate these “all” image pairs, we randomly inpaint five to ten SA-1B masks on image B to simulate change. The text prompt is randomly selected from a predefined list of 96 prompts generated with GPT-4o that convey an intent for all changes.
4.2 Combating Inpainting Noise
Although inpainting produces seemingly realistic changes, we noticed that the inpainted regions tend to have inpainting artifacts (as observed in other works in the literature [9, 19, 10]). We use “red herring” masks to prevent the model from learning artifacts instead of image changes. These red herring masks are a randomly selected subset of zero to ten SA-1B masks that are not classified as any of the classes in the text prompt. Since SA-1B doesn’t provide mask classifications, we use Grounded SAM 2 with a low confidence threshold to ensure that the SA-1B mask doesn’t coincide with a change class instance. These red herring masks are inpainted on image B but not added to the change mask. During training, this forces the model to learn the change between the two images that are consistent with text prompt rather than labeling inpainting noise as changes. Adding red herring masks to “all” image pairs is not applicable, since every change needs to be accounted for which would include all inpaints.
4.3 Dataset Statistics
There is a total of 501,153 change image pairs in CSeg split into 451,090 pairs in train and 50,063 in test. There are 24,298 unique classes and 339,348 unique prompts in train and 7,285 unique classes and 7,326 unique prompts in test. There are 1,408 unique classes and 35,271 unique prompts not seen in train. 53,257 and 5,947 pairs were “all” image pairs in the train and test sets respectively at about a 12% of the total for each set. CSeg’s class counts are top heavy in that more common classes such as “clouds”, “sky”, “people”, “water”, “trees” and “roof” appear exponentially more often than less common classes like “calcite”, “aldravanda”, and “ice hammer” as depicted in Fig. 4. This is due to how the majority of images in SA-1B include these common objects. During class proposals, we select the classes that are least used in CSeg to mitigate this problem.
Although LLaVa-Next and Grounded SAM 2 are very impressive vision models, they cannot produce flawless class proposals and instance masks for all images. This consequently affects the quality of CSeg. To quantitatively measure the quality of CSeg, we manually check randomly sampled images until the percent accuracy clearly converges. We validated 500 images - about 0.1% of the entire dataset - and observed an accuracy of 94.0%. This results in a margin of error of ±2.22% at a 95% statistical confidence. The large majority of errors stem from Grounded SAM 2 yielding masks that are misclassified, especially with more challenging classes. We provide additional examples in Fig. 5 for further evaluation.
Figure 4: Class count distributions: top 1% most frequent classes and the complete dataset. These figures don’t include the “all” class. As class counts decrease, the class gets more impractical.
Figure 5: Evaluation of CSeg quality. Change mask label is highlighted in blue. Manual inspection of the generated images show a 94% agreement with a human reviewer.
5 Experiments
Figure 6: Qualitative Results: We show ViewDelta predictions in red on the test sets, along with the ground truth in blue. All SYSU-CD test images use the text prompt of: “urban development, suburban expansion, pre-construction groundwork, vegetation alteration, road widening, and coastal construction”.
To assess the effectiveness of ViewDelta, we evaluate its performance on three diverse datasets: CSeg, PSCD [36], and SYSU-CD [37]. These datasets cover a broad range of domains, including indoor, outdoor, street-view, and satellite imagery, ensuring a comprehensive evaluation across different contexts.
This section first presents quantitative metrics for each dataset, evaluating several state-of-the-art models on each dataset individually and reporting metrics such as intersection over union (IoU). Following this, we provide a qualitative analysis, showcasing examples of ViewDelta’s performance across various samples in the dataset.
5.1 Quantitative Evaluation
Model IoU F1 Rec Prec
ViewDelta 85.91 92.42 89.25 95.82
Table 1: Quantitative Results On The CSeg Dataset by a jointly trained model, the checkpoint used is the same across all reported results for the general model. The CSeg test set contains prompts not seen in the training set.
Model IoU F1 Rec Prec
Fine-tuned models
CSCDNet + SSCDNet [36] 22.3
CSSCDNet [36] 32.2
ViewDelta 55.54 78.31 69.47 73.47
General model
ViewDelta 52.24 68.63 72.57 65.10
Table 2: Quantitative Results on PSCD Dataset. The highest values are highlighted in bold and the general model results are from the same checkpoint across all results.
Method IoU F1 Rec Prec
Fine-tuned models
FC-EF [11] 61.04 75.81 75.17 76.47
FC-Siam-diff [11] 61.01 75.79 75.30 76.28
FC-Siam-conc [11] 60.23 75.18 76.75 73.67
BiDateNet [30] 62.52 76.94 72.60 81.84
STANet [8] 63.09 77.37 85.33 70.76
DSAMNet [37] 64.18 78.18 81.86 74.81
ChangeFormerV4 [4] 65.03 78.81 77.90 79.74
BIT-50 [6] 66.14 79.62 77.90 81.42
TransUNetCD [23] 66.79 80.09 77.73 82.59
SwinSUNet [49] 68.89 81.58 79.75 83.50
MambaBCD [7] 71.10 83.11 80.31 86.11
ViewDelta 70.09 82.41 79.91 85.08
General model
ViewDelta 68.59 81.37 75.42 88.33
Table 3: Quantitative Results on SYSU-CD Dataset. The highest values are highlighted for First and Second. The general model results are from the same checkpoint across all results. We additionally finetune the general ViewDelta model to evaluate its finetuning performance.
We quantitatively evaluate the performance of ViewDelta on three diverse datasets: CSeg, PSCD [36], and SYSU-CD [37], comparing it against several state-of-the-art methods using standard metrics, including Intersection over Union (IoU), F1 score, Recall (Rec), and Precision (Prec). Results are summarized in Tabs. 1, 2 and 3.
On the CSeg dataset Tab. 1, ViewDelta achieves excellent performance, with an IoU of 85.91%, an F1 score of 92.42%, a recall of 89.25%, and notably high precision of 95.82%.
For the PSCD dataset Tab. 2, which includes viewpoint changes and challenging illumination conditions, ViewDelta significantly outperforms prior fine-tuned methods, achieving an IoU of 55.54%, an F1 score of 78.31%, recall of 69.47%, and precision of 73.47%. The general model also shows strong performance with an IoU of 52.24%.
Finally, on the SYSU-CD dataset Tab. 3, consisting of satellite imagery, the ViewDelta fine-tuned model achieves competitive results with an IoU of 70.09% and an F1 score of 82.41%, closely following the state-of-the-art MambaBCD-Base method. The general model maintains strong performance, achieving an IoU of 68.59% and demonstrating superior precision at 88.33%.
These results demonstrate competitive performance across datasets, underscoring ViewDelta’s robustness. We attribute this success to the model’s multi-domain training approach. Given this strong performance from a relatively simple architecture, we anticipate that further architectural enhancements or specialized fine-tuning approaches could lead to even more substantial performance improvements in the future.
5.2 Qualitative Evaluation
We present qualitative results of our model on example image pairs from the CSeg, PSCD, and SYSU-CD test sets in Fig. 6. For the SYSU-CD test set, all images use the text prompt: “urban development, suburban expansion, pre-construction groundwork, vegetation alteration, road widening, and coastal construction.” For the PSCD test set, each image uses a prompt corresponding to the classes defined in the dataset, such as “vehicle” or “structure.”
5.3 Ablation Study
We perform extensive ablation studies to investigate the impact of various embedding strategies on the training efficiency of ViewDelta. First, we evaluate combinations of image and text embeddings on a random subset of the CSeg dataset containing 22,624 images, allowing models to train for 12 hours. Results are presented in Tab. 4.
Embeddings IoU
SigLip + Dinov2 56.44
SigLip + Patch Embedding 26.16
CLIP + Dinov2 53.46
CLIP + Patch Embedding 27.56
Table 4: Evaluation of embedding combinations on a subset of CSeg.
The results demonstrate that the combination of SigLip text embeddings and Dinov2 image embeddings consistently outperforms other configurations. Additionally, using Dinov2 embeddings significantly surpasses patch embeddings, regardless of text encoder choice.
Further, we evaluate the specific impact of frozen Dinov2 embeddings compared to fully-trained patch embeddings across multiple datasets simultaneously. Both configurations use SigLip for text embeddings, and results are summarized in Tab. 5.
Image Embedding CSeg PSCD SYSU-CD
Frozen Dinov2 85.91 52.24 68.59
Patch Embedding 61.05 16.87 34.81
Table 5: Impact of frozen Dinov2 versus trained patch embeddings across multiple datasets.
We observe that frozen Dinov2 embeddings significantly outperform fully-trained patch embeddings, indicating that pretrained image embeddings not only provide a more robust feature representation but also improve convergence speed and overall model generalization. This advantage is particularly pronounced given the dataset size limitations encountered in our training scenarios.
5.4 Limitations
CSeg is limited to synthetic perspective differences, which limits the models generalizability to large perspective changes. Expanding the dataset to include larger perspective changes is challenging due to the limited availability of the required data in addition to the difficultly in labeling the data. Lack of real data of image pairs with textual prompts. Another limitation is the exponential increase in the number of image tokens required for higher resolution images that may contain small changes. However, this can be mitigated as compute is scaled.
6 Conclusion
We presented ViewDelta, a novel multimodal architecture for text-prompted change detection capable of processing unaligned images through the integration of natural language prompts. By leveraging pretrained text and image embeddings, ViewDelta effectively generalizes across diverse domains without requiring retraining for novel changes. We introduced the CSeg dataset, a comprehensive dataset specifically crafted for text-conditioned change detection tasks, providing over 500,000 image pairs with associated textual prompts and annotated binary masks. Extensive experiments demonstrated that ViewDelta achieves robust and competitive performance across various challenging domains, including indoor, outdoor, street-level, synthetic, and satellite imagery. Our findings underscore the model’s flexibility and effectiveness, paving the way for future advancements in generalized, text-conditioned change detection methodologies.